# LLMTaxo: Leveraging Large Language Models for Constructing Taxonomy of Factual Claims from Social Media

### v1

@@ 3.1.2 Multi-level Topic Generation
To automate the taxonomy construction process and minimize the human effort, we employed LLMs to generate $(t_{b},t_{m},t_{d})$ for each factual claim $c\in\mathcal{C}$. The learning examples $\mathcal{L}$ and the seed taxonomy $\mathcal{S}$ are utilized as part of the prompt for the LLMs. Specifically, the prompt consists of $l_{i}\in\mathcal{L}$, the seed taxonomy $\mathcal{S}$, the instruction and questions that ask LLMs to produce the broad, medium, and detailed topics for $c_{i}$, and the answer to the question (i.e., corresponding $(t_{b_{i}},t_{m_{i}},t_{d_{i}})$ of $c_{i}$). After the LLMs learn from the $k$ examples, it is provided with a new claim $c_{j}$ and asked to generate new topics $(t_{b_{j}},t_{m_{j}},t_{d_{j}})$ for $c_{j}$. Due to the limited context length of the LLMs, one prompt generates $(t_{b_{j}},t_{m_{j}},t_{d_{j}})$ for only one $c_{j}$. This generation process is iterated until finishing generating $(t_{b_{j}},t_{m_{j}},t_{d_{j}})$ for all $c_{j}\in\mathcal{C}$. All the LLM-generated $(t_{b_{j}},t_{m_{j}},t_{d_{j}})$ form $T_{b}$, $T_{m}$, and $T_{d}$. The prompt is detailed in Figure 2 in Appendix A.
3.1.3 Topics Consolidation
After the LLMs generate the topics for each claim, we consolidate the results to build the taxonomy. Medium topics that align with the same broad topic are treated as child nodes of that broad topic, and detailed topics are similarly considered child nodes of their respective medium topics. Formally, we employ two functions: $f_{m}:T_{m}\rightarrow T_{b}$, which maps each medium topic to a broad topic, and $f_{d}:T_{d}\rightarrow T_{m}$, which maps each detailed topic to a medium topic. Therefore, every medium topic is associated with a broad topic: $\forall t_{m}\in T_{m},\quad\exists t_{b}\in T_{b}\text{ such that }f_{m}\left(t_{m}\right)=t_{b}$; similarly, every detailed topic is associated with a medium topic: $\forall t_{d}\in T_{d},\quad\exists t_{m}\in T_{m}\text{ such that }f_{d}\left(t_{d}\right)=t_{m}$.
For instance, consider two claims $c_{1}$ and $c_{2}$ with $\phi(c_{1})$=(Vaccine Safety and Effectiveness, Vaccine aSide Effects, $t_{d}^{(1)}$) and $\phi(c_{2})$=(“Vaccine Safety and Effectiveness,” “Vaccine Injury,” $t_{d}^{(2)}$). By applying $f_{m}$, we get $f_{m}$(“Vaccine Side Effects”)=“Vaccine Safety and Effectiveness” and $f_{m}$(“Vaccine Injury”)=“Vaccine Safety and Effectiveness.” This establishes that “Vaccine Side Effects” and “Vaccine Injury” are child nodes of “Vaccine Safety and Effectiveness.” In this way, we establish a hierarchical structure of topics. This structure ensures that related topics are appropriately grouped, forming the final taxonomy $\mathcal{T}$.
@@ 3.2 Claim Detection
Social media data contains a wide range of content, including personal opinions, personal experiences, and entertainment. Prior to taxonomy construction, we apply claim detection to identify social media posts that are more likely to contain factual claims, which could potentially carry misinformation. We employed a pioneering model, ClaimBuster Hassan et al. (2017), for detecting check-worthy factual claims. This model assigns a score to each sentence or paragraph, indicating the likelihood of it being a check-worthy factual claim. A higher score suggests a greater likelihood of check-worthiness. We set a threshold of 0.5 as it effectively balances precision and recall in identifying check-worthy claims. Posts that scored above this threshold were retained for further steps of $\mathsf{LLMTaxo}$. This approach reduced the dataset to a more manageable subset (see Section 4.1 for details). It also ensures that the retained posts are more likely to contain check-worthy factual information relevant to our taxonomy construction.
@@ 3.3 Identifying Distinct Claims
Many factual claims on social media platforms are frequently repeated or rephrased. For example, the posts “BREAKING: Pentagon rescinds COVID-19 vaccine mandate” and “The Pentagon officially rescinds COVID-19 vaccine mandate” convey the same factual claim but are phrased slightly differently. To reduce redundancy and focus on unique claims, we applied clustering to group identical or nearly identical factual claims. We used HDBSCAN Campello et al. (2013) due to its ability to handle noise and detect outliers, which is particularly useful given that many posts do not closely resemble others and should form individual clusters. To capture the semantic meaning of each claim, we employed Sentence-BERT Reimers (2019) to generate dense vector representations. After clustering, we identified distinct claims by selecting the first post from each cluster while excluding the outlier cluster. Each cluster represents a distinct factual claim, and the identified claims $\mathcal{C}$ form the foundation for taxonomy construction. The outliers represent infrequently discussed content, whereas we only focus on content that appears multiple times. Repeated exposure to information increases belief in its accuracy Pennycook et al. (2018), underscoring the significance of identifying widely circulated claims.
@@ 4.1 Datasets
We collected tweets related to the COVID-19 vaccine using Wildfire Zhang et al. (2024b). For this dataset, we targeted tweets containing various keyword variations related to the COVID-19 vaccine, such as covid19 vaccination, covid-19 vaccine, and covid vax. The data collection period spanned from January 1, 2023 to April 25, 2023, resulting in a total of 384,676 tweets. After applying claim detection aforementioned in Section 3.2, 232,368 tweets were retained for distinct claim identification.
@@
For the climate change dataset, we utilized CrowdTangle CrowdTangle (2024)—a now-discontinued tool as of August 2024—to collect Facebook posts related to climate change. We retrieved posts containing the keyword climate change between January 1, 2024, and May 7, 2024, yielding a total of 229,913 posts. After applying claim detection, we retained 89,412 posts.
@@ 4.2 Implementation Details
For taxonomy construction, we employed two LLMs, Zephyr Tunstall et al. (2023) and GPT-4o mini OpenAI (2024), for performance comparison. Zephyr is selected for its competitive performance in language understanding tasks among all $7$-billion-parameter LLMs Chiang et al. (2024), while GPT-4o mini is chosen for its balance of cost-efficiency and performance. For each dataset, we randomly selected 100 distinct factual claims and annotated them with broad, medium, and detailed topics. From annotated factual claims, we then chose representative samples based on their frequency of occurrence to serve as learning examples for the LLMs. These annotated claims were used to guide the LLMs in generating topics for the distinct claims $\mathcal{C}$ identified through clustering. To minimize the variability in the output, we set the temperature parameter of both models to 0.001. All experiments were conducted on three A100 GPUs.
@@ 4.3.1 Clustering
The statistics of the clusters are shown in Table 1. To evaluate the clusters, we used Silhouette Coefficient Rousseeuw (1987). Since HDBSCAN was employed, the outlier cluster was excluded during the evaluation. The CV dataset achieved the highest silhouette score of 0.940, reflecting highly cohesive and well-separated clusters. Although the CC and CS datasets have lower scores of 0.488 and 0.554, these still indicate reasonably good cluster quality. The variation in silhouette scores of different datasets can be attributed to differences in dataset characteristics, particularly the greater sparsity in Facebook posts, and the Facebook posts are usually lengthier, making it more challenging for clustering. These results suggest that the method performs reasonably well across diverse datasets, even with different clustering configurations.
@@ 4.3.2 Multi-level Topic Generation
The statistics of the generated topics across each dataset are presented in Table 3—the rows labeled “w/” under the method column. It is evident that GPT-4o mini demonstrates a more effective ability than Zephyr in limiting the number of broad and medium topics to a narrower range. However, we observed that some generated topics were associated with only a few factual claims. To enhance the taxonomy’s readability, we consolidated broad topics that appear fewer than 50 times into a new broad category labeled “Other.” For medium and detailed topics, we retained those with occurrences exceeding 4 for medium topics and 3 for detailed topics, respectively. Topics with fewer occurrences were grouped under “Other” topic within their respective parent topics. After merging, the topic statistics of each dataset are shown in Table 2. This approach limits the taxonomy to a more manageable size.
Dataset Model Broad Topic Medium Topic Detailed Topic
CV Zephyr GPT-4o mini 11 8 66 18 114 110
CC Zephyr GPT-4o mini 8 7 146 46 163 229
CS Zephyr GPT-4o mini 10 9 48 25 32 61
Table 2: Topic counts after merging less frequent topics.
4.4 Ablation Study
To evaluate the effectiveness of the presence of the seed taxonomy in the prompt, we conducted experiments with removing the seed taxonomy from the prompt. Specifically, we only provide the LLM with instruction, learning examples, and the target factual claim. The total topic counts with and without seed taxonomy are shown in Table 3. We can see that adding the seed taxonomy to the prompt effectively restricts the taxonomy size, reducing the number of broad topics by up to 99.5%.
Dataset Model Method Broad Topic Medium Topic Detailed Topic
@@ 5 Evaluation
We assessed $\mathsf{LLMTaxo}$ from two aspects: 1) the quality of the taxonomy, and 2) the appropriateness of the generated topics for the factual claims. We engaged both human evaluators and GPT-4 Achiam et al. (2023), applying two sets of evaluation metrics that we have developed to evaluate each aspect. The effectiveness of using LLMs for model performance evaluation has been validated by previous studies Fu et al. (2024); Liu et al. (2023). Identical instructions and metrics were provided to human and GPT-4 evaluators. They were instructed to rate each criterion on a scale from 1 to 5, where 5 indicates the highest quality and 1 the lowest.
To evaluate the quality of taxonomies, we presented taxonomies generated by Zephyr and GPT-4o mini across the three datasets to the evaluators. For the evaluation of the claim-topic suitability, we randomly selected 50 factual claims along with their corresponding lowest-level topics (i.e., leaf nodes of the taxonomy) generated by both models from each dataset, resulting in a total of 100 claim-topic pairs per dataset. We exclusively evaluated the leaf node topics with claims, as the broader taxonomy quality had already been assessed in taxonomy quality evaluation, making further evaluation of higher-level topics unnecessary. Note that not all claims involve three-level topics, as mentioned in Section 3.1.1. Therefore, we focused our evaluation on the leaf node topics. To mitigate bias, we shuffled the claim-topic pairs from both models before presenting them to the evaluators. Both human evaluators and GPT-4 reviewed these pairs using our predefined metrics. Evaluators were also required to provide rationales for their scores. The evaluation prompt is detailed in Appendix C.
@@ 5.1 Taxonomy Evaluation
To design the taxonomy evaluation metrics, we adopted the Goal Question Metric (GQM) approach Caldiera and Rombach (1994) and consulted existing metrics from Kaplan et al. (2022). Due to differences in our tasks, we retained only the orthogonality and completeness metrics from Kaplan et al. (2022). We further refined these metrics to better align with our objectives and introduced additional metrics tailored to our evaluation needs. Each metric is defined with a clear goal, a guiding question, and specified evaluation criteria. An overview of the metrics is provided below, with comprehensive details available in Appendix B.
@@
The goal is to ensure that each topic label communicates its content effectively to avoid confusion. To assess whether the topic labels are clear, precise, and unambiguous, we evaluate precision, unambiguity, consistency, and accessibility.
@@
The goal is to maintain distinct boundaries between topics to ensure that each topic captures unique aspects of the domain. To assess whether the topics are well-differentiated without duplication, we evaluate non-overlap.
@@
The goal is to cover as many areas of the topic to ensure the taxonomy is comprehensive. To assess whether the taxonomy captures a broad and representative set of topics across different aspects of the domain, we evaluate domain coverage, depth, and balance.
For each metric, we calculated the average score of the evaluation criteria for each human evaluator. We then computed the mean of these averages across the three human annotators. The evaluation scores are presented in Table 4. The results affirm the overall efficacy of $\mathsf{LLMTaxo}$, with taxonomies generally receiving high ratings (above 3.3 across all metrics). Notably, GPT-4o mini consistently outperformed Zephyr, suggesting its greater suitability for taxonomy construction. Human evaluators tended to give higher scores than GPT-4, particularly for Zephyr, indicating possibly stricter criteria or different interpretations of taxonomy quality by GPT-4. It is also conceivable that GPT-4 shares more similarities with GPT-4o mini than with Zephyr. The models generally scored well on clarity and completeness, indicating their effectiveness in producing clear, precise and comprehensive topic labels. GPT-4o mini also scores highly in hierarchical coherence, suggesting it does well in structuring information from general to specific in a meaningful way. Orthogonality has slightly lower scores compared to others, particularly for Zephyr, which may indicate some overlap or less distinct boundaries between topics.
@@
CV Human Zephyr 4.2 4.0 3.4 4.3 4.3 4.1
GPT-4o mini 4.3 4.2 3.8 4.7 3.8 3.8
GPT-4 Zephyr 3.5 3.7 3.0 3.7 3.9 4.6
@@
CC Human Zephyr 4.6 4.2 3.9 4.6 3.4 4.6
GPT-4o mini 4.7 4.7 4.2 4.8 4.0 4.0
GPT-4 Zephyr 3.5 3.3 4.0 4.0 3.3 3.4
GPT-4o mini 4.3 4.0 4.5 4.3 3.7 4.1
CS Human Zephyr 4.3 3.9 3.7 4.4 3.9 4.0
GPT-4o mini 4.5 4.0 4.0 4.6 4.4 4.2
GPT-4 Zephyr 3.3 3.0 3.0 3.3 3.8 4.3
GPT-4o mini 4.3 4.7 3.5 4.7 4.1 4.9

### v2

@@ 3.1.2 Multi-level Topic Generation
To automate the taxonomy construction process and minimize the human effort, we employed LLMs to generate topics $(t_{b},t_{m},t_{d})$ for each factual claim $c\in\mathcal{C}$. The learning examples $\mathcal{L}$ and the seed taxonomy $\mathcal{S}$ are utilized as part of the prompt for the LLM. Specifically, the prompt consists of $l_{i}\in\mathcal{L}$ for $i=1,2,\dots,k$ (where $l_{i}=\langle c_{i},(t_{b}^{(i)},t_{m}^{(i)},t_{d}^{(i)})\rangle$ is the claim-topic pair of the $i$-th learning example), the seed taxonomy $\mathcal{S}$, the instruction and questions that ask the LLM to produce the topic tuple for the claim $c$. Due to the limited context length of LLMs, each prompt generates the topic tuple $\phi(c)=(t_{b},t_{m},t_{d})$ for only one $c$. This generation process is iterated until finishing generating topic tuples for all $c\in\mathcal{C}$. The prompt is detailed in Figure 2 in Appendix A.
3.1.3 Topic Consolidation
[…]
For instance, consider two claims $c_{1}$ and $c_{2}$ with the generated topic tuples $\phi(c_{1})$ = (Vaccine Safety and Effectiveness, Vaccine Side Effects, $t_{d}^{(1)}$) and $\phi(c_{2})$ = (“Vaccine Safety and Effectiveness,” “Vaccine Injury,” $t_{d}^{(2)}$). The derived mapping function $f_{m}$ will include $f_{m}$(“Vaccine Side Effects”) = “Vaccine Safety and Effectiveness” and $f_{m}$(“Vaccine Injury”) = “Vaccine Safety and Effectiveness.” This establishes that “Vaccine Side Effects” and “Vaccine Injury” are child nodes of “Vaccine Safety and Effectiveness.”
@@ 3.2 Claim Detection
Social media contains a wide range of content, including personal opinions, personal experiences, and entertainment. Prior to taxonomy construction, we apply claim detection to identify social media posts that are more likely to contain factual claims, which could potentially carry misinformation. We employ the ClaimBuster Hassan et al. (2017) model which assigns a score to each sentence or paragraph, indicating the likelihood of it being a check-worthy factual claim. A higher score suggests a greater likelihood of check-worthiness. We set a threshold of 0.5 as it effectively balances precision and recall in identifying check-worthy claims. Posts that score above this threshold are retained for further steps of $\mathsf{LLMTaxo}$. This approach reduces the dataset to a more manageable subset (see Section 4.1 for details). It also ensures that the retained posts are more likely to contain check-worthy factual information relevant to our taxonomy construction.
@@ 3.3 Identifying Distinct Claims
Many factual claims on social media are frequently repeated or rephrased. For example, the posts “BREAKING: Pentagon rescinds COVID-19 vaccine mandate” and “The Pentagon officially rescinds COVID-19 vaccine mandate” convey the same claim but are phrased slightly differently. To reduce redundancy and focus on unique claims, we apply clustering to group identical or nearly identical factual claims. We use HDBSCAN Campello et al. (2013) due to its ability to handle noise and detect outliers, which is particularly useful given that many posts do not closely resemble others. To capture the semantic meaning of claims, we employ Sentence-BERT Reimers (2019) to generate dense vector representations. After clustering, we identify distinct claims by selecting the first post from each returned cluster list while excluding the outlier cluster. A distinct claim is the representation of a cluster of similar factual claims, and the identified distinct claims $\mathcal{C}$ are used for taxonomy construction. The outliers represent infrequently discussed content, whereas we only focus on content that appears multiple times. Repeated exposure to information increases belief in its accuracy Pennycook et al. (2018), underscoring the significance of identifying widely circulated claims.
@@ 4.1 Datasets
We collected tweets related to COVID-19 vaccines using Wildfire Zhang et al. (2024b). For this dataset, we targeted tweets containing various keyword variations related COVID-19 vaccines, such as “covid19 vaccination,” “covid-19 vaccine,” and “covid vax.” The data collection period spanned from January 1, 2023 to April 25, 2023, resulting in a total of 384,676 tweets. After applying claim detection described in Section 3.2, 232,368 tweets were retained for distinct claim identification.
@@
For the climate change dataset, we utilized CrowdTangle CrowdTangle (2024)—a now-discontinued tool—to collect Facebook posts related to climate change. We retrieved posts containing the keyword “climate change” between January 1, 2024, and May 7, 2024, yielding a total of 229,913 posts. After applying claim detection, we retained 89,412 posts.
@@ 4.2 Implementation Details
For taxonomy construction, we employed three LLMs, Zephyr Tunstall et al. (2023), GPT-4o mini OpenAI (2024), and Gemini 2.0 Flash DeepMind (2025) for performance comparison. Zephyr is selected for its competitive performance in language understanding tasks among all $7$-billion-parameter LLMs Chiang et al. (2024), while GPT-4o mini and Gemini 2.0 Flash are chosen for their balance of cost-efficiency and performance. For each dataset, we randomly selected 100 distinct factual claims and annotated them with broad, medium, and detailed topics. From annotated factual claims, we then chose representative samples based on their frequency of occurrence to serve as learning examples for the LLMs. These annotated claims were used to guide the LLMs in generating topics for the distinct claims $\mathcal{C}$ identified through clustering. All experiments were conducted on three A100 GPUs.
@@ 4.3.1 Clustering
The statistics of the clusters are shown in Table 1. We used Silhouette Coefficient Rousseeuw (1987) to evaluate the clusters’ quality. Since HDBSCAN was employed, the outlier cluster was excluded during the evaluation. The CV dataset achieved the highest Silhouette score of 0.940, reflecting highly cohesive and well-separated clusters. Although the CC and CS datasets have lower scores of 0.488 and 0.554, these still indicate reasonably good cluster quality. The variation in Silhouette scores across different datasets can be attributed to differences in their characteristics, particularly the higher sparsity and greater length of Facebook posts, which make clustering more challenging. These results suggest that the clustering method performs reasonably well across diverse datasets, even with different clustering configurations.
@@ 4.3.2 Multi-level Topic Generation
The statistics of the generated topics across datasets are presented in Table 2—the rows labeled “w/” under the method column. It is evident that Gemini and GPT-4o mini outperformed Zephyr in constraining the number of broad and medium topics to a narrower range. Gemini also demonstrated superior ability in limiting the number of detailed topics. However, we observed that some generated topics were associated with only a few factual claims. To enhance the taxonomy’s readability, we consolidated broad topics that appear fewer than 50 times into a new broad category labeled “Other.” For medium and detailed topics, we retained those with occurrences exceeding 4 for medium topics and 3 for detailed topics, respectively. Topics with fewer occurrences were grouped under “Other” topic within their respective parent topics. After merging, the topic statistics of each dataset are shown in Table 3. This approach limits the taxonomy to a more manageable size.
Dataset Model Method
Broad
Topic
Medium
Topic
Detailed
Topic
@@
Dataset Model
Broad
Topic
Medium
Topic
Detailed
Topic
CV
Zephyr
GPT-4o mini
Gemini
11
8
9
66
18
27
114
110
12
$C$
Zephyr
GPT-4o mini
Gemini
8
7
7
146
46
22
163
229
30
CS
Zephyr
GPT-4o mini
Gemini
10
9
9
48
25
32
32
61
26
Table 3: Topic counts after merging less frequent topics.
4.4 Ablation Study
To evaluate the effectiveness of the presence of the seed taxonomy in the prompt, we conducted experiments with removing the seed taxonomy from the prompt. Specifically, we only provide the LLMs with instructions, learning examples, and the target factual claims. The total topic counts with and without seed taxonomy are shown in Table 2. We can see that adding the seed taxonomy to the prompt effectively restricts the taxonomy size, reducing the number of broad topics by up to 99.5%.
@@ 5 Evaluation
We assessed $\mathsf{LLMTaxo}$ from two perspectives: 1) the quality of the generated taxonomy, and 2) the relevance of the generated topics to the factual claims. We developed a set of evaluation metrics to assess these two aspects. We engaged both human evaluators and GPT-4 Achiam et al. (2023) for the evaluation. The effectiveness of using LLMs for model performance evaluation has been validated by previous studies Fu et al. (2024); Liu et al. (2023). Identical instructions and metrics were provided to both human and GPT-4 evaluators. They were instructed to rate each criterion on a scale from 1 (strongly disagree) to 5 (strongly agree), where 5 indicates the best quality and 1 the worst.
To evaluate the quality of taxonomies, we presented taxonomies generated by Zephyr, GPT-4o mini, and Gemini across the three datasets to the evaluators, resulting in a total of 9 taxonomies. For the evaluation of the claim-topic suitability, we randomly selected 102 factual claims along with their corresponding lowest-level topics (i.e., leaf nodes of the taxonomy) generated by the three models from each dataset, resulting a total of 306 claim-topic pairs for evaluation. We exclusively evaluated the leaf node topics with claims, as the broader topics had already been assessed in taxonomy quality evaluation. Note that not every claim has three-level topics, as mentioned in Section 3.1. Therefore, we focused our evaluation on the leaf node topics. To mitigate bias, we shuffled claim-topic pairs from different models before presenting them to the evaluators. Both human evaluators and GPT-4 reviewed these pairs using our predefined metrics. Human evaluators were required to provide rationales for the scores. The evaluation prompt for GPT-4 is detailed in Appendix B.
@@ 5.1 Taxonomy Evaluation
To design the taxonomy evaluation metrics, we adopted the Goal Question Metric (GQM) approach Caldiera and Rombach (1994) and consulted existing metrics from Kaplan et al. (2022). Due to differences in our tasks, we retained only the orthogonality and completeness metrics from Kaplan et al. (2022). We further refined these metrics to better align with our objectives and introduced additional metrics tailored to our evaluation needs. Each metric is defined with a clear goal, a guiding question, and specified evaluation criteria. An overview of the metrics is provided below, with comprehensive details available in Appendix C.
@@
The goal of this metric is to ensure that each topic label communicates its content effectively to avoid confusion. To assess whether the topic labels are clear, precise, and unambiguous, we evaluate precision, unambiguity, consistency, and accessibility.
@@
The goal is to maintain clear boundaries between topics, ensuring that each one captures a unique aspect of the overall topic domain. To assess whether the topics are well-differentiated and free from redundancy, we evaluate their distinctiveness and degree of overlap.
@@
The goal of this metric is to cover as many areas of the topic domain as possible to ensure the taxonomy is comprehensive. To assess whether the taxonomy captures a broad and representative set of topics across different facets of the topic domain, we evaluate its domain coverage, depth, as well as balance.
Three human evaluators and GPT-4 were provided with evaluation instructions and the background of the taxonomy construction. Each of them rated the taxonomies based on the evaluation criteria. For each metric, we calculated the average score of the evaluation criteria for each human evaluator. We then computed the mean of these averages across the three human annotators. The evaluation scores are presented in Table 4.
The results affirm the overall efficacy of $\mathsf{LLMTaxo}$, with taxonomies generally receiving high ratings (above 3.0 across all metrics). Notably, GPT-4o mini consistently outperformed Zephyr and Gemini, suggesting its greater suitability for taxonomy construction. Human evaluators tended to give higher scores than GPT-4, particularly for taxonomies generated by Zephyr and Gemini, indicating possibly stricter criteria or different interpretations of taxonomy quality by GPT-4. It is also reasonable to assume that GPT-4 shares greater similarity with GPT-4o mini than with Zephyr or Gemini.
The models generally scored well on clarity and completeness, indicating their effectiveness in producing clear, precise and comprehensive topic labels. GPT-4o mini also scored higher in hierarchical coherence with the highest score of 4.7, suggesting it does well in structuring information from general to specific in a meaningful way. Orthogonality has slightly lower scores compared to others, particularly for Zephyr and Gemini, which indicates some overlap or less distinct boundaries between topics. We observed Gemini was more strongly influenced by the seed taxonomy and tended to generate topics closely aligned with it. This is reflected in the limited number of unique topics produced by Gemini, as shown in Table 3 and Table 2. Such alignment negatively affected both orthogonality and completeness. To assess the agreement level among human evaluators, we computed the inter-rater reliability using Gwet’s AC2 Gwet (2008) with quadratic weights, which yielded a score of 0.80.
@@
CV Human Zephyr 4.2 4.0 3.3 4.3 4.7 3.6
GPT-4o mini 4.3 4.2 3.8 4.7 4.4 4.6
Gemini 4.4 3.9 3.5 3.9 4.4 4.6
GPT-4 Zephyr 3.5 3.7 3.0 3.7 4.0 4.6
@@
Gemini 3.4 3.0 3.0 3.3 4.0 4.4
$C$ Human Zephyr 4.6 4.2 3.9 4.6 3.3 3.9
GPT-4o mini 4.7 4.7 4.2 4.8 4.4 4.5
Gemini 4.3 3.9 3.3 4.7 4.4 4.6
GPT-4 Zephyr 3.5 3.3 4.0 4.0 3.4 3.4
GPT-4o mini 4.3 4.0 4.5 4.3 3.8 4.5
Gemini 3.8 3.7 3.5 3.7 3.8 4.4
CS Human Zephyr 4.3 3.9 3.7 4.4 3.8 4.0
GPT-4o mini 4.5 4.0 4.0 4.6 4.7 4.6
Gemini 4.6 3.6 3.5 3.9 4.8 4.6
GPT-4 Zephyr 3.3 3.0 3.0 3.3 3.7 4.0
GPT-4o mini 4.3 4.7 3.5 4.7 4.0 4.9
Gemini 3.0 3.0 3.5 3.3 4.0 4.9